### v1

# Analytic Personalized Federated Meta-Learning

Abstract
Analytic federated learning (AFL) which updates model weights only once by using closed-form least-square (LS) solutions can reduce abundant training time in gradient-free federated learning (FL). The current AFL framework cannot support deep neural network (DNN) training, which hinders its implementation on complex machine learning tasks. Meanwhile, it overlooks the heterogeneous data distribution problem that restricts the single global model from performing well on each client’s task. To overcome the first challenge, we propose an AFL framework, namely FedACnnL, in which we resort to a novel local analytic learning method (ACnnL) and model the training of each layer as a distributed LS problem. For the second challenge, we propose an analytic personalized federated meta-learning framework, namely pFedACnnL, which is inherited from FedACnnL. In pFedACnnL, clients with similar data distribution share a common robust global model for fast adapting it to local tasks in an analytic manner. FedACnnL is theoretically proven to require significantly shorter training time than the conventional zeroth-order (i.e. gradient-free) FL frameworks on DNN training while the reduction ratio is $98\%$ in the experiment. Meanwhile, pFedACnnL achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings, compared with the previous SOTA frameworks.
Index Terms:
gradient-free, analytic federated learning, meta-learning.
I Introduction
The rapid expansion of internet-of-things (IoT) devices, increasing from under 9 billion in 2019 to over 15 billion by 2023 [1], has been a key driver in the development of Federated Learning (FL). FL is a privacy-preserving technique of Distributed Collaborative Machine Learning (DCML) that enables model training across multiple edge devices without uploading raw data to a central parameter server (PS). Training a general FL model contains three typical steps: (i) a set of edge devices conduct local computation for model training, usually by the stochastic gradient descent (SGD) method, based on their own dataset, and upload the resultant model parameters to the server, (ii) the server aggregates the uploaded parameters to generate a new global model, and (iii) the server feeds back the new global model to edge devices for the next round of local model training. This procedure repeats until the loss function converges and a certain model training accuracy is achieved.
Typical FL training schemes include first- [2, 3] (e.g., SGD) and second-order methods [4], which necessitate the computation of gradients and Hessian matrices. Despite the efficiency of these gradient descent (GD)-based methods, their reliance on gradient computations can pose challenges in practice. Specifically, (i) edge devices with poor computational capabilities are incapable of performing frequent gradient calculations. For instance, mobile phones or IoT sensors are primarily designed for inference-only purposes rather than backpropagation [5]. In such cases, only small, pre-trained or pruned models are computationally affordable for execution on these devices [1, 6]. (ii) Gradient information is not always available to support gradient computations. For instance, in situations such as federated hyperparameter tuning [7] and federated black-box attack [8], the gradients cannot be derived due to the non-differentiable nature of the loss function. As a result, GD-based FL frameworks may become infeasible in these scenarios, prompting a strong demand from both academia and industry for the development of gradient-free FL training approaches.
Recently, the zeroth-order FL framework has been widely adopted as an efficient gradient-free FL training approach [5, 9, 10]. The conventional zeroth-order FL frameworks, including BAFFLE [5] and FedZO [9], conduct the finite difference technique to approximate gradients, effectively substituting the backpropagation in gradient computations with multiple inference processes. However, such gradient approximation is always biased, necessitating intensive sampling processes to reduce the approximation error. Consequently, the training time required by these frameworks is excessively long. To solve this problem, analytic federated learning (AFL) [11] has been proposed as an effective and emerging approach without any gradient approximation or computations.
In AFL, the model weights are updated only once and are determined by using closed-form least-square (LS) solutions. Therefore, AFL can achieve a significantly shorter training time than the conventional zeroth-order FL frameworks. However, AFL overlooks the performance limitations of training a single global model for clients with heterogeneous data distributions. In other words, training a unified global model to satisfy the performance requirements of all clients is challenging due to the distinct preferences in data generation or collection by different clients. Personalized FL (PFL) [12] is a promising paradigm to alleviate the heterogeneous data distribution problem. In PFL, each client utilizes the common knowledge, which is the aggregated global model on the PS, to generate a personalized model tailored for its local data distribution. Meta-learning [13, 14, 15], as a representative model optimization technique in PFL, enables the aggregated global model (i.e. the meta-model) to fast adapt to heterogeneous tasks by performing several gradient descents on the global model. This motivates us to introduce meta-learning to AFL to address the heterogeneous data distribution problem in AFL. This introduction results in two challenges: First of all, the current AFL framework [11] does not have the capability of training deep neural networks (DNN), which is the basis of tackling complex machine learning tasks. Thus, this nature makes the introduction of meta-learning to the current AFL framework meaningless when we confront complex machine learning tasks. Secondly, the existing meta-learning method is only suitable for clients using GD-based optimization methods. Therefore, combining meta-learning with AFL directly is impractical since AFL does not involve any gradient approximation or computations.
In this paper, we first resort to ACnnL (analytic convolutional neural network learning) [16], an analytic single-device learning method for DNN, and propose its federated implementation, called FedACnnL, to enable DNN training for AFL. Specifically, for the PS, a layer-wise global model update rule using closed-form LS solutions is designed with a specialized aggregation strategy to realize the layer-wise collaborative DNN training. For the clients, an adaptive batch size algorithm is incorporated to combat the straggler problem of FL, in which the PS must keep idle until all clients finish their local computations. By increasing the local batch size to reduce the local computation time of lagging clients, the idle time of the PS is diminished. Thanks to the weight-invariant property of FedACnnL, this incorporation achieves acceleration without any model performance degradation. Therefore, FedACnnL, as an AFL framework, realizes computation-efficient collaborative DNN training. Then, based on FedACnnL, we propose an analytic personalized federated meta-learning (APFML) framework, called pFedACnnL. A gradient-free meta-learning method is designed in it to combat the heterogeneous data distribution problem. Specifically, the PS groups the clients with similar data distribution to generate a meta-model that contains common knowledge within each group by following the standard training process of FedACnnL. Next, each client can achieve its personalized model tailored for its local data distribution without using any gradient information by analytically solving a local PFL objective. This objective bridges the gap between the personal knowledge of its local dataset and the common knowledge of its meta-model. Our contributions can be summarized as follows:
  • We first resort to ACnnL and propose FedACnnL to enable computation-efficient DNN training in AFL. With a competitive model performance on DNN, FedACnnL is proven to require a greatly shorter training time than the conventional zeroth-order (i.e. gradient-free) FL frameworks both theoretically and empirically.
  • Based on FedACnnL, we propose pFedACnnL, which is the first APFML framework to the best of our knowledge. A gradient-free meta-learning method is designed in it for fast adaption of the group-wise meta-models to local tasks, thereby promoting the model performance on each client under the heterogeneous data distribution setting.
  • We validate the efficiency of our frameworks on a CPU-cluster-based testbed, which can emulate the environmental settings of data heterogeneity and resource heterogeneity in FL simultaneously. Experimentally, FedACnnL can reduce the training time by $88\%\sim 98\%$ compared with the conventional first- and zeroth-order FL frameworks. Meanwhile, pFedACnnL achieves state-of-the-art (SOTA) model performance in most cases of convex and non-convex settings, compared with other previous SOTA first- and zeroth-order PFL frameworks.
TABLE I: Table of partial notations.
  Notations Description
  $n$ the index of a sample
$i$ the index of a client
$b$ the index of a local batch
$\textbf{x}_{n}$ the $n$-th data vector of the entire dataset $\mathcal{S}$
X the data matrix of the entire dataset $\mathcal{S}$
$\textbf{X}^{(i)}$ the data matrix of the local dataset on the $i$-th client
$\textbf{X}_{l}^{[n]}$ the real input matrix of the $l$-th layer using $\textbf{x}_{n}$ as DNN input
$\textbf{X}_{l}^{(i)}$ the real input matrix of the $l$-th layer
on the $i$-th client using $\textbf{X}^{(i)}$ as DNN input
$\textbf{X}^{\{b\}}$ the $b$-th batch of data matrix
$\textbf{X}_{l}^{\{b\}}$ the real input matrix of the $l$-th layer using $\textbf{X}^{\{b\}}$ as DNN input
$\textbf{y}_{n}$ the $n$-th label vector of the entire dataset $\mathcal{S}$
Y the label matrix of the entire dataset $\mathcal{S}$
$\textbf{Y}^{(i)}$ the label matrix of the local dataset on the $i$-th client
$\bar{\textbf{Z}}_{l}^{[n]}$ the pseudo label matrix of the $l$-th layer by encoding $\textbf{y}_{n}$
$\bar{\textbf{Z}}_{l}^{(i)}$ the pseudo label matrix of the $l$-th layer
on the $i$-th client by encoding $\textbf{Y}^{(i)}$
$\textbf{Y}^{\{b\}}$ the $b$-th batch of label matrix
$\bar{\textbf{Z}}_{l}^{\{b\}}$ the pseudo label matrix of the $l$-th layer by encoding $\textbf{Y}^{\{b\}}$
$(\cdot)^{T}$ the transpose of a matrix
[…]
II Preliminaries
II-A Notations
[…]
II-B Analytic Convolutional Neural Network Learning (ACnnL)
Unlike the prior analytic learning methods [17, 18, 19] that only support shallow DNN training or training the single output layer with a pre-trained backbone, ACnnL enables DNN training. AcnnL projects the label information into the hidden layers so that each hidden layer is equipped with supervised information (i.e. pseudo labels). Then, it updates the model weights layer-wisely by using closed-form LS solutions in a supervised learning style with the layer input and pseudo labels. Since the real label lies within the pseudo one’s linear space, mapping the input features onto the pseudo labels is almost equivalent to mapping it onto the real label, thereby achieving DNN training [16]. Meanwhile, thanks to the layer-wise weight update mechanism, ACnnL completes model training in only one epoch, in which the number of iterations matches that of the DNN layers. In the following, we recap the training process in each iteration.
Figure 1: The resampling and flattening process in convolutional layers to make the weights in it updatable by the equation 2.
Given an $N$-sample dataset $\mathcal{S}=\{\textbf{x}_{n},\textbf{y}_{n}\}_{n=1}^{N}$ where $\textbf{x}_{n}\in{\mathbb{R}}^{d_{0}}$ and $\textbf{y}_{n}\in{\mathbb{R}}^{d_{y}}$ denote the data vector and the label vector of the $i$-th sample respectively. We stack these samples to generate the data matrix $\textbf{X}=\left[\textbf{x}_{1};\textbf{x}_{2};\cdots;\textbf{x}_{N}\right]\in\mathbb{R}^{N\times{d_{0}}}$ and the label matrix $\textbf{Y}=\left[\textbf{y}_{1};\textbf{y}_{2};\cdots;\textbf{y}_{N}\right]\in\mathbb{R}^{N\times{d_{y}}}$. In the $l$-th iteration, the pseudo label matrix of the $l$-th hidden layer $\bar{\textbf{Z}}_{l}$ is generated as $\bar{\textbf{Z}}_{l}=\textbf{YQ}_{l}$, where ${\textbf{Q}_{l}}\in{\mathbb{R}^{d_{y}\times{d_{l}}}}$ is the label encoding matrix which is randomly initialized from a normal distribution. The encoding is linear so that $\bar{\textbf{Z}}_{l}$ can fit the dimensionality of $\textbf{Z}_{l}\in\mathbb{R}^{N\times{d_{l}}}$, which is the real output matrix before entering the activation function of the $l$-th hidden layer $f_{l}(\cdot)$ when X is the input of the DNN model being trained. Specially, $\bar{\textbf{Z}}_{l}=\textbf{Y}$ when training the output layer of the model. Then, ACnnL formulates the supervised learning of the $l$-th layer as an least-square (LS) problem with $l_{2}$ regularization:
$\underset{\textbf{W}_{l}}{\arg\min}~{}\|{\bar{\textbf{Z}}_{l}-\textbf{X}_{l}\textbf{W}_{l}}\|_{2}^{2}+\gamma\|{\textbf{W}_{l}}\|_{2}^{2}$ (1)
where $\textbf{X}_{l}\in\mathbb{R}^{N\times{d_{l-1}}}$ and $\textbf{W}_{l}\in\mathbb{R}^{d_{l-1}\times{d_{l}}}$ denote the real input matrix and model weights of the $l$-th hidden layer respectively. Specially, $_{l}=\textbf{X}$ X when training the input layer. $\gamma$ is the regularization coefficient which is a hyperparameter to control the search space of the LS solutions. Next, we can take the closed-from LS solutions of the weights in $l$-th hidden layer by putting the derivatives with respect to $\textbf{W}_{l}$ on the objective function in the equation 1 to 0. We present the solutions at a sample level as follows:
$\hat{\textbf{W}}_{l}=(\sum_{n=1}^{N}{\textbf{X}_{l}^{[n]T}\textbf{X}_{l}^{[n]}}+\gamma{\textbf{I}})^{-1}(\sum_{n=1}^{N}\textbf{X}_{l}^{[n]T}\bar{\textbf{Z}}_{l}^{[n]})$ (2)
where $\hat{\textbf{W}}_{l}$ denotes the trained model weights of the $l$-th layer. Specially, when ACnnL trains convolutional layers, $\textbf{X}_{l}^{[n]}\in\mathbb{R}^{W_{l-1}\times{H_{l-1}}\times{C_{l-1}}}$ is an image and $\textbf{W}_{l}~{}(i.e.~{}{\mathcal{W}_{l}})=\{\textbf{W}_{l}^{[c]}\}_{c=1}^{C_{l}},\textbf{W}_{l}^{[c]}\in\mathbb{R}^{ks\times{ks}\times{C_{l-1}}}$ is the set of convolution kernels, where $W,H,C$ are width, height and channel respectively, and $ks$ is the kernel size. As illustrated in Fig. 1, ACnnL reformulates the supervised learning of convolutional layers to be compatible with the equation 1 by resampling on the $\textbf{X}_{l}^{[n]}$ with a sliding window of $ks$ size and flattening the $\mathcal{W}_{l}$ to $\textbf{W}_{l}\in\mathbb{R}^{(ks\times{ks}\times{C_{l-1}})\times{C_{l}}}$. After the flattened $\textbf{W}_{l}$ is replaced by $\hat{\textbf{W}}_{l}$ generated from the equation 2, $\hat{\textbf{W}}_{l}$ is transformed into $\hat{\mathcal{W}}_{l}$, which has the original shape of $\mathcal{W}_{l}$.
After ACnnL completes the aforementioned label encoding stage and the weight update stage, it computes the real input matrix of $(l+1)$-th layer, $\textbf{X}_{l+1}=f_{l}(\textbf{X}_{l}\hat{\textbf{W}}_{l})~{}or~{}f_{l}(\textbf{X}_{l}{\star}\hat{\mathcal{W}}_{l})$ (where $\star$ is the convolution operation) and finally steps into the next iteration to train the next layer. ACnnL finishes its one-epoch model training when all layers of DNN are trained. Fig. 2 pictures the workflow in the arbitrary iteration of ACnnL.
Figure 2: The workflow of ACnnL in the arbitrary iteration. The red dashed box represents the current training layer while the blue dashed box indicates the next layer to be trained.
III Design of FedACnnL
In this section, we first model the weight update problem in each layer of ACnnL as a distributed LS problem. Then, based on the distributed LS problem, we propose the federated implementation of ACnnL, called FedACnnL, followed by the adaptive batch size algorithm incorporated in it. Finally, we give a theoretical analysis on the complexity of the computation and communication in FedACnnL.
III-A Distributed LS Problem
Based on the supervised learning property of ACnnL in each hidden layer, we can model the weight update problem in each layer as a distributed LS problem. Specifically, suppose we have the aforementioned $N$-sample dataset $\mathcal{S}$ in the section II-B, and $C$ clients of edge devices. The dataset is disjointly distributed onto the $C$ clients, i.e. $\mathcal{S}=\cup_{i=1}^{C}\mathcal{D}_{i},\ {\ \mathcal{D}}_{m}\cap\mathcal{D}_{n}=\emptyset,\ \forall\ m\neq n$, where $i$ is the index of a client and $\mathcal{D}_{i}$ denotes the local dataset of a client. Each client stacks the local dataset to generate the local data matrix $\textbf{X}^{(i)}$ and the local label matrix $\textbf{Y}^{(i)}$ like section II-B. Then, in the arbitrary $l$-th hidden layer, each client encodes $\textbf{Y}^{(i)}$ with a same global $\textbf{Q}_{l}$ to generate a local pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}\textbf{Q}_{l}$. Specially, $\bar{\textbf{Z}}_{l}^{(i)}=\textbf{Y}^{(i)}$ when the $l$-th hidden layer is the output layer. With the real input matrix at the $l$-th layer $\textbf{X}_{l}^{(i)}$ and the pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}$, we can formulate the distributed/federated optimization problem at the arbitrary $l$-th layer as:
$\begin{split}&\underset{\textbf{W}_{l}}{\arg\min}~{}\sum_{i=1}^{C}F(\textbf{W}_{l};\mathcal{D}_{i})\\ where~{}F(\textbf{W}_{l}&;\mathcal{D}_{i})=\|{\bar{\textbf{Z}}_{l}^{(i)}-\textbf{X}_{l}^{(i)}\textbf{W}_{l}}\|_{2}^{2}+\gamma\|{\textbf{W}_{l}}\|_{2}^{2}\\ \end{split}$ (3)
Assume that the model weights prior to the $l$-th layer on all clients are identical, in other words, all $\textbf{X}^{(i)}$ undergo the same computational operations to obtain $\textbf{X}_{l}^{(i)}$. Then, we can get the following two facts:
$\begin{split}&\sum_{i=1}^{C}{\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}=\sum_{n=1}^{N}{\textbf{X}_{l}^{[n]T}\textbf{X}_{l}^{[n]}}\\ &\sum_{i=1}^{C}\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}=\sum_{n=1}^{N}\textbf{X}_{l}^{[n]T}\bar{\textbf{Z}}_{l}^{[n]}\\ \end{split}$ (4)
That is because the overall dataset formed by all local datasets of the clients is equivalent to the original $N$-sample dataset and the sum of all intermediate output matrices is not changed. Finally, the federated optimization problem can be analytically solved using the equation 2 and the equation 4:
$\hat{\textbf{W}}_{l}=(\underbrace{\sum_{i=1}^{C}{\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}}_{\textbf{A}_{0}}+\gamma{\textbf{I}})^{-1}(\underbrace{\sum_{i=1}^{C}\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}}_{\textbf{A}_{1}})$ (5)
For convenience, we define the sum results of the two facts as $\textbf{A}_{0}$ and $\textbf{A}_{1}$ respectively in the rest of the paper. Based on the distributed LS problem, we propose the federated implementation of ACnnL, i.e. FedACnnL, in the next subsection.
III-B Federated Implementation of ACnnL
Before we step into the description of the training rounds with a same workflow, we first present the initialization at the beginning of FedACnnL: Firstly, each client initializes a local DNN model with the same structure. Then, the PS broadcasts the same random seed $s$ to all clients. Finally, each client utilizes $s$ to generate an identical set of label encoding matrices, i.e. $\mathcal{Q}=\{\textbf{Q}_{l}\}_{l=1}^{L-1}$, where $L$ is the total number of the DNN layers being trained.
Based on the distributed LS problem defined above, we give a detailed description of the workflow in the arbitrary $l$-th training round, in which clients collaboratively train the $l$-th hidden layer of the DNN model, as follows:
Local Computation. Firstly, each client allocates the memory space for $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$, by creating zero matrices with the same shapes of them. Then, each client randomly samples a fixed batch of data matrix i.e. $\textbf{X}^{\{b\}}\in\mathbb{R}^{B_{l}^{(i)}\times{d_{0}}}$ and label matrix $\textbf{Y}^{\{b\}}\in\mathbb{R}^{B_{l}^{(i)}\times{d_{y}}}$, where $b$ is the index of the local batch and $B_{l}^{(i)}$ is the local batch size of the $i$-th client in the $l$-th training round, from the local dataset $\mathcal{D}_{i}$ without replacement until no batch of data can be generated from it. In each sampling process, each client utilizes the previously trained layers, i.e. $\{\hat{\textbf{W}}_{1},\cdots,\hat{\textbf{W}}_{l-1}\}$, which are identical over all clients, to generate a batch of real input matrix $\textbf{X}_{l}^{\{b\}}$ by the forward pass of $\textbf{X}^{\{b\}}$ from the input layer to the $(l-1)$-th layer. Meanwhile, the pseudo label matrix $\bar{\textbf{Z}}_{l}^{\{b\}}$ is obtained by the label encoding technique, i.e. $\bar{\textbf{Z}}_{l}^{\{b\}}=\textbf{Y}^{\{b\}}\textbf{Q}_{l}$ (or $\bar{\textbf{Z}}_{l}^{\{b\}}=\textbf{Y}^{\{b\}}$ when the $l$-th layer is the output layer). The $\textbf{X}_{l}^{\{b\}T}\textbf{X}_{l}^{\{b\}}$ and $\textbf{X}_{l}^{\{b\}T}\bar{\textbf{Z}}_{l}^{\{b\}}$ is added to the $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$ respectively at the end of each sampling process. To summarize, the local computation process of the $l$-th training round can be mathematically represented as:
$\begin{split}&\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}=\sum_{b=1}^{\lceil|\mathcal{D}_{i}|/B_{l}^{(i)}\rceil}{\textbf{X}_{l}^{\{b\}T}\textbf{X}_{l}^{\{b\}}}\\ &\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}=\sum_{b=1}^{\lceil|\mathcal{D}_{i}|/B_{l}^{(i)}\rceil}\textbf{X}_{l}^{\{b\}T}\bar{\textbf{Z}}_{l}^{\{b\}}\\ \end{split}$ (6)
where $|\mathcal{D}_{i}|$ is the total number of samples in the local dataset of the $i$-th client. There exists a fact in the equation 6 that regardless of what number $B_{l}^{(i)}$ is, the resultant $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$ remain invariant. This fact indicates the weight-invariant property of AFL, which is further discussed in the latter of this subsection.
Aggregation. The resultant $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$ are uploaded to the PS by each client. After the PS receives the resultant two items from all clients, it aggregates the items respectively by a sum operation, i.e. $\sum_{i=1}^{C}{\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}$ and $\sum_{i=1}^{C}\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$, to generate the $\textbf{A}_{0}$ and $\textbf{A}_{1}$ in the equation 5. Then, the PS utilizes the $\textbf{A}_{0}$ and $\textbf{A}_{1}$ to update the global model weights of the $l$-th layer by using the closed-from LS solutions in the equation 5. Finally, the PS feeds the trained global model weights at the $l$-th layer, i.e. $\hat{\textbf{W}}_{l}$, back to all clients.
Local Weight Update. After clients receive the $\hat{\textbf{W}}_{l}$ from the PS, they replace the local model weights of the $l$-th layer with the global one, i.e. $\hat{\textbf{W}}_{l}$. Then, they step into the next training round to collaboratively train the next layer, i.e. $(l+1)$-th layer until all layers of their local DNN model have been trained.
Weight-Invariant Property. Based on the facts in the equation 4 and 6, we can find that the $\textbf{A}_{0}$ and $\textbf{A}_{1}$ in the equation 5 remain invariant unless the overall dataset formed by all local datasets of the clients, or the set of label encoding matrices $\mathcal{Q}$ changes. Therefore, no matter how the overall dataset is distributed among the clients, no matter what number the batch size is, and no matter how the model weights are initialized on the clients, the final closed-form LS solutions of each layer will not be changed in FedACnnL.
III-C Adaptive Batch Size Algorithm
In the real-world FL scenario, the computing power and the network conditions of the clients are various. Thus, the overall time consumption of the clients completing the local computation and uploading the aforementioned resultant two items are inconsistent. This inconsistency is the straggler problem of FL, in which the PS must keep idle until the resultant items of the client with the poorest computing power and network condition arrive at the PS. An empirical practice to combat this problem is to increase the local batch size of the lagging clients to fully exert the parallelism of the hardware (CPUs or GPUs), thereby reducing the local computation time of the lagging clients. As a result, the item arrival time of the lagging clients at the PS can be advanced, reducing the idle duration of the PS. In our design, the PS adaptively adjusts the local batch size for each client in each training round by computing the ratio of each client’s time consumption to the global minimum one. The adjusted result is then sent back to each client with the trained global model weights obtained in the aggregation step. This adjustment can be mathematically represented as:
$B_{l+1}^{(i)}=min(\lceil\frac{T_{l}^{(i)}-T_{l-1}^{(i)}}{min(\{T_{l}^{(i)}-T_{l-1}^{(i)}\}_{i=1}^{C})}\times{B_{l}^{(i)}}\rceil,B_{max}^{(i)})$ (7)
where $T_{l}^{(i)}$ denotes the item arrival time of the $i$-th client at the PS in the $l$-th training round and $T_{0}^{(i)}=0$. The time interval between two consecutive item arrivals of the $i$-th client, i.e. $T_{l}^{(i)}-T_{l-1}^{(i)}$ is an approximation of the client’s overall time consumption (including computation and communication) in the $l$-th training round. $B_{max}^{(i)}$ is the largest local batch size permitted to the $i$-th client and is equal to $|\mathcal{D}_{i}|$. Notably, the local batch size of each client is adjusted in every training round due to the dynamic network condition. Last but not least, the adjustment of the local batch size for each client cannot change the final model weights of each layer on each client, thanks to the weight-invariant property. Therefore, the introduction of the adaptive batch size algorithm does not influence the final model performance on each client. The pseudo-code of FedACnnL with the adaptive batch size algorithm is shown in the algorithm 1.
Input : The participating clients, the regularization coefficient $\gamma$, the initial local batch size $B_{1}^{(i)}$, the random seed $s$;
Output : A global trained model $\{\hat{\textbf{W}}_{l}\}_{l=1}^{L}$;
1 Client initialization: initialize a local DNN model $\{\textbf{W}_{l}\}_{l=1}^{L}$ and generate a same set of label encoding matrices $\mathcal{Q}=\{\textbf{Q}_{l}\}_{l=1}^{L-1}$ by $s$;
2 for $l=1~{}to~{}L$ do
3       for Clients $i=1~{}to~{}C$ in parallel do
4            Initialize $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$ respectively with zero matrices;
5             for batch of samples $\textbf{X}^{\{b\}},\textbf{Y}^{\{b\}}$} with the local batch size $B_{l}^{(i)}$ sampled from $\mathcal{D}_{i}$ without replacement do
6                   $\textbf{X}_{1}^{\{b\}}=\textbf{X}^{\{b\}}$; $\bar{\textbf{Z}}_{l}^{\{b\}}=\textbf{Y}^{\{b\}}$;
7                   if $l>1$ then
8                        $\textbf{X}_{l}^{\{b\}}={f_{l-1}(\underbrace{\cdots{f_{1}(\textbf{X}_{1}^{\{b\}}\hat{\textbf{W}}_{1})\cdots}}_{\textbf{X}_{l-1}^{\{b\}}}\hat{\textbf{W}}_{l-1})}$;
9                         if $l<L$ then
10                               $\bar{\textbf{Z}}_{l}^{\{b\}}=\bar{\textbf{Z}}_{l}^{\{b\}}\textbf{Q}_{l}$;
11                         end if
12
13                   end if
14                  $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}+=\textbf{X}_{l}^{\{b\}T}\textbf{X}_{l}^{\{b\}}$;
15                   $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}+=\textbf{X}_{l}^{\{b\}T}\bar{\textbf{Z}}_{l}^{\{b\}}$;
16
17             end for
18            Upload $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$ to the PS;
19
20       end for
21      The PS does:
22       Record the item arrival time $T_{l}^{(i)}$ of each client;
23       Generate $B_{l+1}^{(i)}$ using the equation 7 for each client;
24       $\textbf{A}_{0}=\sum_{i=1}^{C}{\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}$; $\textbf{A}_{1}=\sum_{i=1}^{C}\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$;
25       Generate $\hat{\textbf{W}}_{l}$ using the equation 5;
26       Send $\hat{\textbf{W}}_{l}$ and $B_{l+1}^{(i)}$ back to each client;
27       Each client does:
28       Replace $\textbf{W}_{l}$ with $\hat{\textbf{W}}_{l}$;
29       Set the local batch size to $B_{l+1}^{(i)}$;
30 end for
Algorithm 1 FedACnnL with the adaptive batch size algorithm
III-D Theoratical Analysis on Computation and Communication
To prove the superiority of our proposed gradient-free framework (i.e. FedACnnL) on the training time, we compare the computational and communication complexity of it with other gradient-free frameworks (i.e. the conventional zeroth-order frameworks: FedZO [9] and BAFFLE [5]) theoretically. For a fair comparison, we assume that the local number of samples and the local batch size on each client are equal and fixed (i.e. $|\mathcal{D}_{i}|=|\mathcal{D}|$, $B_{l}^{(i)}=B$) during the whole FL process of each framework.
Computational complexity. In the conventional paradigm of FL, the whole training process contains a number of training rounds and each training round requires each client to conduct several repeated visits (i.e. local epochs) on its whole local dataset. Meanwhile, each epoch includes multiple iterations, in which each client samples a batch of data and labels from $\mathcal{D}_{i}$ without replacement and conducts certain computations on the batch of data and labels. Since each DNN layer in FedACnnL converges in one training round which contains one local epoch, FedACnnL only requires $L$ training rounds in total and each training round contains one local epoch. That is greatly smaller than both FedZO and BAFFLE, in which more than 20/40 training rounds (each contains more than one local epoch) are required to achieve a certain model performance empirically on the MNIST [20]/CIFAR-10 [21] dataset using LeNet-5 (5 layers) [22]. Therefore, our analysis focuses on the computations conducted on the batch of data and labels to prove a lower batch-level computational complexity for FedACnnL. In FedACnnL, the batch-level computations contain one forward pass and three matrix multiplications (line 8 and line 10 $\sim$ 14 in the algorithm 1). However, in FedZO and BAFFLE, the batch-level computations incorporate performing a large number of finite difference processes to accurately approximate the gradient. In each finite difference process, two forward passes should be conducted. After removing the computational complexity of one forward pass for each framework, the complexity of the remaining three matrix multiplications in FedACnnL is $\mathcal{O}(B(d_{y}d_{l}+d_{l-1}^{2}+d_{l-1}d_{l}))$ in the arbitrary $l$-th training round. Meanwhile, more than one forward passes exists in FedZO or BAFFLE and each of them has a complexity of $\mathcal{O}(\sum_{l=1}^{L}Bd_{l-1}d_{l})$. Thus, even if FedZO or BAFFLE merely conducts the finite difference process once (which is impractical) in the batch-level computations, FedACnnL can still save a considerable amount of training time. This theoretical conclusion is empirically verified in the section V-B.
Communication complexity. In each training round, each client in FedACnnL uploads the resultant two items (i.e. $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$) to the PS and receive the trained global model weights of the current training layer $\hat{\textbf{W}}_{l}$ from the PS. Notably, $B_{l}^{(i)}$ is too small and we ignore it. Thus, the total communication complexity for FedACnnL in the arbitrary $l$-th training round is $\mathcal{O}(d_{l-1}^{2}+d_{l-1}d_{l}+\theta_{l})$, where $\theta_{l}$ denotes the parameter number of the $l$-th DNN layer. Meanwhile, each client in FedZO or BAFFLE uploads and receives the entire DNN model in each training round, resulting in a communication complexity of $\mathcal{O}(2\sum_{l=1}^{L}\theta_{l})$ per round. Whether in the convolutional layer or the fully-connected layer, $\theta_{l}$ is equal to $d_{l-1}d_{l}$. Therefore, FedACnnL can achieve a lower communication complexity than FedZO and BAFFLE unless the output dimensionality of a certain layer exceeds the parameter number of the entire DNN model. That is a rare case. Furthermore, since FedACnnL requires only $L$ training rounds while FedZO and BAFFLE need a great number one empirically (see section V-B), the total communication complexity of FedACnnL in the whole FL process is much lower than FedZO and BAFFLE.
IV Design of pFedACnnL
IV-A Overview of pFedACnnL
To tackle the heterogeneous data distribution problem that restricts the single global model from performing well on each client’s task, we propose pFedACnnL, a personalized version of FedACnnL. pFedACnnL encompasses three stages: the initialization stage, the federated optimization stage and the local personalization stage. An overview of pFedACnnL is illustrated in Fig. 3.
Figure 3: The overview of pFedACnnL. The cyan dashed box represents the initialization stage while the brown dashed box and the orange dashed box represent the federated optimization stage and the local personalization stage respectively.
The initialization stage. In the same way as FedACnnL, each client utilizes the random seed $s$ received from the PS to generate an identical set of label encoding matrices $\mathcal{Q}$. Then, a local model is initialized with the same structure on each client. Next, different from FedACnnL, pFedACnnL requires each client to upload an encoded label vector, which is the representation of the local data distribution on each client, to the PS. Finally, the PS uses these encoded label vectors to conduct the $K$-means clustering algorithm and classifies the clients into $K$ groups. In this way, pFedACnnL enables the clients that have similar local data distribution to train a joint meta-model, namely the group-wise meta-model, in the next stage. By doing so, each group-wise meta-model can contain a specialized unique common knowledge within its group and make it easier for its clients to conduct local personalization than a single global model (see section V-C).
The federated optimization stage. In each group, the clients follow the standard process of FedACnnL. Specifically, in the arbitrary $l$-th federated training round, each client conducts the local computation and uploads the resultant two items (i.e. $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$) to the PS as in FedACnnL. Next, the PS aggregates the items respectively by a sum operation within each group and then generates the model weights of the $l$-th layer for the $K$ group-wise meta-models respectively. At the end of the federated training round, each client receives the model weights belonging to its group to update the $l$-th layer of the local model. After pFedACnnL completes its all federated training rounds, $K$ group-wise meta-models are born across the clients.
The local personalization Stage. Each client uses the local meta-model to achieve a personalized model catering for the local data distribution by using a gradient-free meta-learning method. This method follows a similar optimization process to ACnnL but updates the model weights of each layer using closed-from LS solutions of a local PFL objective. Specifically, in the arbitrary $l$-th local personalization iteration, each client generates its pseudo label matrix and real input matrix of the $l$-th hidden layer as in ACnnL. Then, a local PFL objective is designed to achieve supervised learning of the $l$-th hidden layer and meta-learning from the group-wise meta-model simultaneously. The personalized model weights of the $l$-th hidden layer are obtained by analytically solving the objective on each client. After that, each client computes the real input matrix of the next layer and steps into personalizing the next layer. The local personalization stage finishes when all clients have personalized each layer of their local meta-models.
IV-B The Initialization Stage
After each client initializes the local model and the set of label encoding matrices $\mathcal{Q}$ in the same way as FedACnnL, it uploads an encoded label vector to the PS for conducting the $K$-means clustering algorithm. Specifically, the encoded label vector indicates the information of the local data distribution. This information can be directly represented by the sum of all local label vectors, i.e. $\sum_{\textbf{y}_{n}\in{\mathcal{D}_{i}}}\textbf{y}_{n}$ for the arbitrary $i$-th client. However, the sum of the label vectors can reveal the client’s preference or privacy. Therefore, to safeguard the privacy to some extent and reduce the communication burden, the sum of the label vectors should be normalized and then encoded by $\textbf{Q}_{enc}$ which is selected from $\mathcal{Q}$ and satisfies the following equation:
$\textbf{Q}_{enc}=\underset{\textbf{Q}_{l}\in{\mathcal{Q}}}{\arg\min}~{}d_{l}$ (8)
Then, the uploaded encoded vector for the arbitrary $i$-th client (i.e. $\textbf{h}_{i}$) can be mathematically represented as:
$\textbf{h}_{i}=\textbf{Q}_{enc}\frac{1}{|\mathcal{D}_{i}|}\sum_{\textbf{y}_{n}\in{\mathcal{D}_{i}}}\textbf{y}_{n}=\frac{1}{|\mathcal{D}_{i}|}\sum_{\textbf{y}_{n}\in{\mathcal{D}_{i}}}\textbf{Q}_{enc}\textbf{y}_{n}$ (9)
After the PS receives the encoded vectors $\{\textbf{h}_{i}\}_{i=1}^{C}$ from all clients, it utilizes these vectors to conduct the $K$-means clustering algorithm to classify the clients into $K$ groups, i.e. $\{\mathfrak{S}_{k}\}_{k=1}^{K}$, where $k$ is the index of a group and $\mathfrak{S}_{k}$ denotes a sequence of clients’ indices. As a result, the clients belonging to the same group have similar local data distribution, resulting in a potential common knowledge within each group to be tapped in the following federated optimization stage.
IV-C The Federated Optimization Stage
In this stage, the clients belonging to the same group collaboratively train a joint meta-model by following the standard process of FedACnnL. Specifically, in the arbitrary $l$-th federated training round of the $k$-th group, pFedACnnL reformulates the federated optimization problem in the $l$-th layer of the group-wise meta-model by reconstructing the distributed LS problem in the equation 3 as follows.
$\begin{split}&\underset{\textbf{M}_{l}^{[k]}}{\arg\min}~{}\sum_{i\in{\mathfrak{S}_{k}}}F(\textbf{M}_{l}^{[k]};\mathcal{D}_{i})\\ where~{}F(\textbf{M}_{l}^{[k]}&;\mathcal{D}_{i})=\|{\bar{\textbf{Z}}_{l}^{(i)}-\textbf{X}_{l}^{(i)}\textbf{M}_{l}^{[k]}}\|_{2}^{2}+\gamma\|{\textbf{M}_{l}^{[k]}}\|_{2}^{2}\\ \end{split}$ (10)
where $\textbf{M}_{l}^{[k]}$ denotes the meta-model weights of the $l$-th layer in the $k$-th group. Then, we can present the closed-form LS solutions of the equation 10 by putting its derivatives with respect to $\textbf{M}_{l}^{[k]}$ to 0:
$\hat{\textbf{M}}_{l}^{[k]}=(\underbrace{\sum_{i\in{\mathfrak{S}_{k}}}{\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}}_{\textbf{A}_{0}^{[k]}}+\gamma{\textbf{I}})^{-1}(\underbrace{\sum_{i\in{\mathfrak{S}_{k}}}\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}}_{\textbf{A}_{1}^{[k]}})$ (11)
where $\hat{\textbf{M}}_{l}^{[k]}$ denotes the trained meta-model weights of the $l$-th layer in the $k$-th group. $\textbf{A}_{0}^{[k]}$ and $\textbf{A}_{1}^{[k]}$ are the aggregated results of the resultant two items (i.e. $\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}$ and $\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}$) uploaded by the clients in the $k$-th group. Based on the reformulation, we re-describe the federated optimization stage elaborately. In the arbitrary $l$-th federated training round, each client conducts its local computation as in FedACnnL and uploads the resultant two items to the PS. Then, the PS generates the aggregated results for each group respectively, i.e. $\{\textbf{A}_{0}^{[k]},\textbf{A}_{1}^{[k]}\}_{k=1}^{K}$ and utilizes them to obtain $\{\hat{\textbf{M}}_{l}^{[k]}\}_{k=1}^{K}$ for each group, by using the equation 11. At the end of the federated training round, each client receives the trained meta-model weights of the $l$-th layer belonging to its group and uses them to replace the local model weights of the $l$-th layer. The federated optimization stage completes when all layers of the local model on each client have been replaced once. By the way, the adaptive batch size algorithm is also applied in pFedACnnL. Specifically, before the PS aggregates the resultant two items in each federated training round, it records the item arrival times for all clients and calculates the local batch size of the next round for each client by using the equation 7. At the end of each federated training round, the PS sends the calculated result accompanied with the trained meta-model weights back to each client.
IV-D The Local Personalization Stage
After the federated optimization stage, each client obtains a meta-model that incorporates the common knowledge applicable to all clients’ tasks within its group. The aim of the local personalization stage on each client is to generate a personalized model from its meta-model to promote the model performance on each client’s task. For a fast generation in our gradient-free setting, a gradient-free meta-learning method is proposed and incorporated in this stage. Specifically, each client personalizes and updates the local model weights in a layer-wise way like ACnnL. In the arbitrary $l$-th local personalization iteration of the $i$-th client, the client gets its pseudo label matrix $\bar{\textbf{Z}}_{l}^{(i)}$ and real input matrix $\textbf{X}_{l}^{(i)}$ as in ACnnL. Then, pFedACnnL formulates the personalization of the $l$-th layer on the $i$-th client as an LS problem with the following PFL objective:
$\underset{\textbf{W}_{l}^{(i)}}{\arg\min}~{}\underbrace{\|{\bar{\textbf{Z}}_{l}^{(i)}-\textbf{X}_{l}^{(i)}\textbf{W}_{l}^{(i)}}\|_{2}^{2}}_{g(\cdot)}+\epsilon\underbrace{\|{\textbf{W}_{l}^{(i)}-\hat{\textbf{M}}_{l}^{[k]}}\|_{2}^{2}}_{h(\cdot)}$ (12)
where $\textbf{W}_{l}^{(i)}$ denotes the model weights of the $l$-th layer on the $i$-th client. $g(\cdot)$ denotes the supervised learning of the $l$-th hidden layer while $h(\cdot)$ denotes the meta-learning from the client’s meta-model. $\epsilon$ is the hyperparameter that controls the tradeoff between the personalization performance and the generalization performance in FL. By putting the derivatives of the equation 12 with respect to $\textbf{W}_{l}^{(i)}$ to 0, we can get the personalized model weights of the $l$-th layer on the $i$-th client, i.e. $\hat{\textbf{W}}_{l}^{(i)}$, analytically as:
$\hat{\textbf{W}}_{l}^{(i)}=({\textbf{X}_{l}^{(i)T}\textbf{X}_{l}^{(i)}}+\epsilon{\textbf{I}})^{-1}(\textbf{X}_{l}^{(i)T}\bar{\textbf{Z}}_{l}^{(i)}+\epsilon{\hat{\textbf{M}}_{l}^{[k]}})$ (13)
After replacing the local model weights with the personalized ones, each client computes the real input matrix of the next layer, i.e. $\textbf{X}_{l+1}^{(i)}=f_{l}(\textbf{X}_{l}^{(i)}\hat{\textbf{W}}_{l}^{(i)})~{}or~{}f_{l}(\textbf{X}_{l}^{(i)}{\star}\hat{\mathcal{W}}_{l}^{(i)})$, and steps into the next local personalization iteration. The local personalization stage finishes when all layers of the local meta-model have been personalized.
V Performance Evaluation
V-A Experimental Settings
Testbed. We implement a prototype of our proposed frameworks on a CPU-cluster-based testbed with 101 nodes. The first node is treated as the PS while the other nodes is regarded as the clients. Each node is equipped with four Intel(R) Xeon(R) Platinum 9242 CPUs. We implement one process on each node with a restriction on the available number of CPU cores. The available number of CPU cores on each node is set randomly ranging from 1 to 4 to emulate the edge devices that have different computing power and network conditions. The nodes are connected by InfiniBand network with a bandwidth of 100Gbps.
Datasets. We consider classification tasks using three datasets: the MNIST dataset [20], the synthetic dataset [23] and the CIFAR-10 dataset [21] to simulate three distinct levels of task difficulty in FL (from easy to hard). The MNIST dataset is a handwritten digit dataset with 70,000 instances, in which each sample contains a binary image with a resolution of $28\times{28}\times 1$ and a label indicator ranging from 0 to 9. The synthetic dataset is produced artificially following the methodology outlined in [23], integrating two parameters, $\bar{\alpha}$ and $\bar{\beta}$, to govern the degree of distinction between the local model and dataset for every individual client. The CIFAR-10 dataset contains 60,000 tiny RGB images, which have a resolution of $32\times{32}\times 3$ and can be classified into 10 categories. For the MNIST and CIFAR-10 datasets, we employ the Dirichlet sampling $p_{u,i}{\sim}Dir_{N}(\beta)$ to generate the heterogeneous sample partitions of the datasets over 100 clients, where $p_{u,i}$ denotes the probability of allocating a sample of label $u$ to the $i$-th client. A smaller $\beta$ means more diversity in the data distribution between local clients. We set $\beta=0.1$ to emulate the highly heterogeneous data environment [24]. For the synthetic dataset, we adhere the power law distribution to distribute the samples among 100 clients. Meanwhile, we set $\bar{\alpha}$ and $\bar{\beta}$ as 0.5 as in [25].
Baselines and Metircs. To prove the superiority of our proposed AFL (i.e. gradient-free FL) framework, FedACnnL, on the efficiency of training DNN, we compare it with the following four frameworks (including GD-based and gradient-free frameworks) and choose the averaged total training time of each client as the main metric. The total training time of each client consists of the computation time, the idle time and the computation time. The computation time is the time cost for conducting local computation. The idle time for a client is the time span during which it awaits the completion of the local computation by the client with the poorest computing power. The communication time is the time expense of transmitting data between the client and the PS.
  • FedACnnL-WA: This is the version of FedACnnL without the adaptive batch size algorithm. The existence of it is to evaluate the effectiveness of the introduced adaptive batch size algorithm to our AFL framework.
  • FedAvg [2]: It is the classical first-order FL framework that enables multiple clients collaboratively to train a high-quality global model.
  • FedZO [9]: It is the standard zeroth-order FL framework that enables each client to approximate the local gradients with only DNN inference processes by using the finite difference technique.
  • BAFFLE [5]: A novel baseline for zeroth-order FL frameworks, which resorts to the finite difference technique to approximate the gradients locally on each client.
Furthermore, to evaluate the effectiveness of our proposed APFML framework, pFedACnnL (the personalized version of FedACnnL), we compare it with the vanilla FedACnnL and another four PFL frameworks. We choose the overall test accuracy of all clients as the metric of the model performance.
  • FedACnnL: The vanilla version of our proposed AFL framework, which is directly used to prove the benefits of its personalized version on the model performance.
  • pFedACnnL-WC: This is the version of pFedACnnL without the $K$-means clustering algorithm. In other words, the federated optimization stage completely follows the standard process of FedACnnL without any in-group aggregation. The existence of it is to prove the necessity of the component (i.e. the $K$-means clustering algorithm) in pFedACnnL.
  • pFedACnnL-WP: This is the version of pFedACnnL without the local personalization stage. After the federated optimization stage, the framework kills the process. The existence of it is to prove the necessity of the component (i.e. the local personalization stage) in pFedACnnL.
  • pFedMe [15]: A first-order PFL framework that utilizes Moreau envelops, which is a meta-learning technique to find a balance between generalization and personalization performance in FL.
  • pFedZO [25]: A zeroth-order PFL framework that extends pFedMe to FedZO by replacing the backpropagation for gradient computation with the finite difference technique for gradient approximation.
[…]
Hyperparameter setting. We choose $\gamma=100$, $\epsilon=2500$ and $K=10$ for FedACnnL and pFedACnnL since this combination achieves the best model performance empirically in our experiments. In our first experiment, in which we compare FedACnnL with other frameworks in the averaged total training time, the number of training rounds is set to 20 on the MNIST dataset according to [5], while it is set to 600 on the synthetic dataset according to [15]. In the second experiment, in which we evaluate our APFML framework, pFedACnnL and its variants choose the test accuracy after they finish their all stages as the final model performance. Meanwhile, all other frameworks choose the best test accuracy within 1,000 training rounds as the final model performance, just as in [25].
Figure 4: The averaged total training time of each client in each framework on the MNIST and synthetic datasets.
Figure 5: The overall test accuracy of all clients versus training round cure during the training procedure on the MNIST and synthetic datasets.
V-B Training Efficiency
To highlight the efficiency of FedACnnL on training DNN, we carry out the experiment in the non-convex setting (i.e. DNN) to validate the training time reduction of FedACnnL over the aforementioned first- and zeroth-order FL frameworks (i.e. FedAvg, FedACnnL-WA, FedZO and BAFFLE). Before we discuss the training time, we should establish the model performance guarantee for FedACnnL. Fig. 5 demonstrates the overall test accuracy of all clients versus training round curve during the training procedure. It shows that FedACnnL can approach a competitive model performance with the first-order FL framework FedAvg on the MNIST dataset while it achieves the state-of-the-art (SOTA) model performance on the synthetic dataset. Meanwhile, FedACnnL-WA achieves the same model performance as FedACnnL due to the weight-invariant property which has been discussed in the section III-B. On the basis of the performance superiority, we present the averaged total training time for each framework in Fig. 4. As illustrated in Fig. 4, FedACnnL can reduce 88%, 98% and 99% training time of FedAvg, FedZO and BAFFLE respectively while the reduction ratio is more than 98% on the synthetic dataset. That is because FedACnnL updates its model weights only once in a layer-wise manner while the other frameworks require clients to iteratively update their models locally and need substantial training rounds. Furthermore, the total time per training round in FedZO and BAFFLE is $4.7s,7.0s$ on the MNIST dataset and $0.31s,1.07s$ on the synthetic dataset while that in FedACnnL is $0.42s$ and $0.15s$ on the MNIST and synthetic datasets. This fact validates our theoretical analysis in the section III-D which indicates that FedACnnL has a lower complexity of computation and communication on each client in each training round than FedZO and BAFFLE. Finally, FedACnnL reduces $13\%$ and $21\%$ idle time of FedACnnL-WA on the MNIST and synthetic dataset, proving the effectiveness of the adaptive batch size algorithm in which the time expenses of the local computation for the lagging clients are reduced.
Figure 6: The model performance and training time of our proposed frameworks on the CIFAR-10 dataset using the DCNN model.
Performance on a deeper and wider DNN model. To further prove the effectiveness of our proposed frameworks on the DNN training, we implement FedACnnL, pFedACnnL and their variants on training the aforementioned DCNN model. Fig 6 shows the model performance and training time of our proposed frameworks on the CIFAR-10 dataset using the DCNN model. Since pFedACnnL-WP, FedACnnL-WA and FedACnnL do not have the local personalization stage, we fill in the test accuracy of the blank training rounds (i.e. rounds from 5 to 9), which are used for local personalization in pFedACnnL and pFedACnnL-WC, for them with their final test accuracy. FedACnnL can approach a competitive model performance with FedAvg (70.06% versus 73%) while its personalized version i.e. pFedACnnL-WP and pFedACnnL can achieve a test accuracy of 74% and 74.40%, which is 1% better than FedAvg. Furthermore, all of our proposed frameworks can finish their training in a minute-level total training time.
TABLE II: Table of model performance.
  Settings Dataset-Model Frameworks
FedACnnL pFedACnnL-WC pFedACnnL-WP pFedMe pFedZO pFedACnnL
[…]
Synthetic-DNN $86.09(\pm 0.06)$ $88.63(\pm 0.01)$ $89.22(\pm 0.03)$ $88.58(\pm 0.17)$ $89.05(\pm 0.04)$ $\textbf{90.00}(\pm 0.03)$
convex MNIST-LR $87.11(\pm 0.00)$ $\textbf{96.24}(\pm 0.00)$ $91.79(\pm 0.01)$ $93.22(\pm 0.03)$ $93.09(\pm 0.05)$ $95.81(\pm 0.01)$
Synthetic-LR $81.80(\pm 0.00)$ $87.81(\pm 0.00)$ $88.33(\pm 0.01)$ $83.13(\pm 0.01)$ $72.14(\pm 0.08)$ $\textbf{88.96}(\pm 0.01)$
V-C Personalized Model Performance
We evaluate our proposed APFML framework in both the convex and non-convex settings. The evaluation is done by taking the mean and standard deviation of the test accuracy from running the experiment for 5 times. Table II lists the model performance of pFedACnnL, the vanilla FedACnnL and another four PFL frameworks. In the table, we can find that pFedACnnL experiences a performance improvement of $4\%\sim 6\%$ in the test accuracy under the non-convex setting while it has an improvement of $7\%\sim 8\%$ under the convex setting, compared to the vanilla FedACnnL. Furthermore, as illustrated in Fig. 6 and Table II, in most cases (except in the MNIST dataset), pFedACnnL has a better model than its versions without a certain component (i.e. pFedACnnL-WC and pFedACnnL-WP). For instance, in Fig. 6, pFedACnnL achieves a test accuracy of 74.40% on the CIFAR-10 dataset while pFedACnnL-WC and pFedACnnL-WP achieve a test accuracy of 72.88% and 74.03% respectively. Notably, on the CIFAR-10 and synthetic datasets, compared with pFedACnnL-WP, pFedACnnL-WC which uses a single global model has an even lower model performance than pFedACnnL, proving the effectiveness of the grouping mechanism in the highly heterogeneous data environment. The special case is that pFedACnnL-WC outperforms pFedACnnL in the MNIST dataset. That is because the task difficulty of the MNIST dataset is comparatively lower than that of the synthetic and CIFAR-10 datasets. Thus, introducing too many complex techniques (including DNN, group-wise meta-learning and etc) to pFedACnnL for easy machine learning tasks can cause an over-fitting problem. Last but not least, pFedACnnL achieves the SOTA performance compared to the first- and zeroth-order PFL frameworks (i.e. pFedMe and pFedZO) in most cases while it achieves a competitive performance with pFedMe (96.23% versus 96.92%) in the non-convex setting of the MNIST dataset.
VI Conclusions
In this paper, we have proposed FedACnnL to enable computation-efficient DNN training in AFL and proposed pFedACnnL based on it to mitigate the heterogeneous data distribution problem in AFL. Specifically, FedACnnL is the federated implementation of ACnnL with an introduction of the adaptive batch size algorithm to boost the training procedure while pFedACnnL incorporates a gradient-free meta-learning method on the basis of FedACnnL. Experimentally, FedACnnL has a significantly shorter training time than the conventional first- and zeroth-order FL frameworks while pFedACnnL achieves the SOTA model performance in most cases of convex and non-conv setting.